Wednesday, July 27, 2011

Life - Mu Alpha Theta 2011 Review, Pt 1

• So hello to all. I'm still a bit jet lagged, but I thought I'd just post my review of Mu Alpha Theta 2011 in Dallas.


•Overall, it was... an okay experience. I have to be honest, I enjoyed some parts and hated others, but overall I think my favorite trip was last year's in D.C. Anyway, allow me to review different aspects of the convention.


•This review will have 3 parts. Part 1, the hotel and location; Part 2, the math tests; and Part 3, other convention events not covered in the first two parts.
•So, the hotel...


Rooms: 7/10
•The suites were very nice, but our room smelled a lot like cigarette smoke. Ew. The beds were comfy enough and there was a good amount of space, but I thought the lights could've been brighter. Nothing much to complain about, but nothing to rave about either. The rooms were clean and comfortable, though a bit cold, and that was good enough.
Service: 2/10
•Absolutely HORRIFYING experience with the service. First, the lock on the door to my room broke, and both times the front desk assumed it was me and my roommates' faults for leaving our cards next to our phones and demagnetizing them. NO. Our door was broken. But they didn't believe us. And on a related note, because the door was broken the maid couldn't get in, so she just skipped over our room and NEVER REPORTED IT.
•Also, the maids failed to leave enough towels for the girls in our room (5 of us, but 4 towels?).
•When our team arrived, they also told us only one room was ready, so we had to squeeze our whole team in there for THREE HOURS before the rest were finally ready to go. Urgh.
•I thought that the ants in my room at the 2009 convention were bad, but this just takes the pi. >: ( We were there for a math convention. Our team needed time and space to study and do our events, but the bad service definitely made it harder for us to do that. FAIL.
Food: 9/10
• This was some of the best convention food I ever had! The selection for breakfast was great and I loved the little fruit tarts and other desserts we had on the last night. 
The food was not perfect, but considering the convention food of the past, this was excellent. Hands down this was the best out of all 3 years.
Overall Impression: 5/10
• The hotel was okay. The food may have been good, but the sucky service makes me hope they NEVER hold another convention there again. Honestly, as long as the food is edible and no one's getting sick, that's fine for a MAO convention. But it makes it waaay too hard on the students when the service is crappy and people are getting locked out of their rooms. Like I said before, we needed time to study, and some of that time was taken away thanks to the idiots who work at the hotel.


• Annnd that's all I've got for now. I'll post my next two portions of the review later.

Saturday, June 25, 2011

Mathy Facts - Codes and Ciphers

• I was going to call this "Fun Facts", but I guess some people might not find these facts all that fun, so whatever...
• Anyway, as I slightly somewhat sort-of mentioned in an earlier post, one of my events at Mu Alpha Theta this summer is an open topic event on Codes and Ciphers. So here's ten fun facts I learned about the topic from the two books I've been reading on the subject, Codes, Ciphers and Secret Writing by Martin Gardner and Cracking Codes and Cryptograms for Dummies by Denise Sutherland and Mark E. Koltko-Rivera.


1. A code is defined as being a kind of "new language", while a cipher is a message with letters that are replaced or rearranged by other symbols. (Ciphers tend to be easier to crack than codes, and most people mistake ciphers as codes.)
2. Plaintext is the secret message and ciphertext is the encrypted message.
3. The most often used letters, in order, are: E, T, A, O, and N.
4. The most frequently used two letter word is OF.
5. The most frequently used three letter word is THE.
6. In English, the letter Q must always be followed by U.
7. The most common double letters, in order, are: LL, EE, SS, OO, and TT.
8. Julius Caesar used two ciphering systems when he was general of the Roman armies, known as "Caesar Shift" and "Caesar Box Codes".
9. The Freemasons use codes to keep information on Masonic ceremonies secret.
10. A transposition cipher is a cipher where the letters in the message are the same, but their positions are changed.


• Isn't it interesting? :D ...no?


• Well, I'll post more interesting information later. I'm getting started on studying integrals next so I'm sure something interesting will turn up there! ;D


Less Than Three,
The Artist

Wednesday, June 15, 2011

Formula Fun - Sequences and Series

• So today I was at math practice and we were studying sequences and series!!! ^o^ And a teammate of mine showed me a very important formula to remember:


∞ _ _ _ _ _ _ _ _ _  _
∑ (x^n) / (n!) = e^x
n=0 _ _ _ _ _ _ _ _ _ _ _ _  _

• This is the Taylor series expansion for e^x. It's very important to know in sequences and series, because Taylor series is good for approximating the value of a function (at a certain value). He had to have this memorized for when he took the BC AP Calculus exam because he is a ninja when it comes to math (and life). I'm not as smart, however, so I only took AB and didn't learn it until today. :c Still, I'm very grateful he pointed this formula out to me. And hopefully posting it on my blog will help me to remember it! :D Isn't it just so awesome and amazing and interesting and useful?!?! yeah, not really unless you're studying sequences i guess...

• But anyway, I hope you enjoy this wonderful formula. c: Until next time!

Less than three,
The Artist

Monday, May 23, 2011

Life - Summer is almost here!

• So... our team didn't do so well at states... not even in the trophy range. TT__TT However, all of us had a lot of fun anyway... or at least I did. I'm kind of the least serious out of all the teammates in my grade, I think. Sometimes I feel like some of them should lighten up more.


BUT ANYWAY


• I'm excited because summer is nearly here! And that means time to focus on Mu Alpha Theta! ^_^ Oh, and no school. That too.


I'm a Mu this year, so I signed up for Integration (M), Codes and Ciphers (O), and Sequences and Series (M). I'm also going to be working our poster this year! ^_^ I wanted to sign up for Quizzle too, but it overlaps with the time to work on the poster, so Coach said I can't. >_<


• Well, it's really late and I should be getting to bed. With the extra time from summer, hopefully I can post my first comic up soon! 


I'll share anything interesting I learn too. ^_^


Less than three,
The Artist

Monday, April 18, 2011

Probability Lesson - Dice Games, Part 1/3

• Introduction: One thing type of math that I have a definite love/hate relationship with is probability. On the one hand, I find it amusing that school is actually teaching the basics of good gambling. On the other hand, it makes me pull out my hair in frustration. Well, at least it's not as bad as calculus; I curse the freaking math gods for ever inventing that torture.


Math team season is now over in my state, and other than the seniors' only state event coming up next month and Mu Alpha Theta nationals during the summer, competition for this year is over. But nationals are important and I still need to prepare, so I figured: hey! Why not actually put the blog I made to good use? Coach always says that teaching others helps to reinforce the ideas, so this should help me out, right?


Thus, this first post is born.


Oh, and for anyone reading an planning on subscribing because you think I'm going to teach stuff that'll help you do your homework: find another website. I can guarantee that my posts will be 50% math comics or something related, 25% funny math stories, 20% math that's kind of awesome but not really anything to help you in class, and 5% anything that will actually have relevance in pre-calculus or whatever the hell you're taking.


First mini-lesson in probability: You've got a 1/20 chance that this will be helpful for class, unless your teacher is actually doing a lesson focused on dice games and how to design them so you can take people's money. If you do have a teacher like that, then I just have to say you're one lucky S.O.B.


So come along with me on this magical journey as I teach you all how to figure out your odds at a basic dice game. I'll cover a couple of examples and show you how to design one that's in your favor.


Let's begin with an example, shall we?


• O • O • O • O • O •

• Example 1: Your friend invents a dice game for you to play. He has two standard, 6-sided dice. Both are fair and have equal odds of rolling on any of the 6 sides. You roll both dice and add the numbers. If the total is a multiple of 4, you win $8. If not, you pay your friend $3. Is it a good idea to play this game?


How does this deal sound right about now? It certainly seems like the odds could be in your favor, right? After all, $8 is over twice as much as $3. But don't just be stupid and just jump in. A good gambler needs to be a good mathematician.


The first step is to dissect problem to get the information we want. What are you looking for? Well, to figure out if this is a good idea or not, we need to figure out whether you or your friend has the higher chance of making money. In other words, we're looking for the expected win/loss per play.


5 Steps to find "expected win/loss per play" (for this problem):


• 1. First, figure out all of the possible combinations to get a winning number, in this case, a multiple of four. Since the maximum you can roll is 12 (6+6), that means the only winning rolls will have to add up to 4, 8, and 12.


• 2. Next, let's find all the possible combinations. (NOTE: Remember to treat the each dice roll as separate events, since the outcome of one roll doesn't directly depend on the other)


If your first roll is a 1, your next roll would need to be a 3 to equal 4.
If your first roll is a 2, your next roll would need to be a 2 to equal 4, or a 6 to equal 8.
If your first roll is a 3, your next roll would need to be a 1 to equal 4, or a 5 to equal 8.
If your first roll is a 4, your next roll would need to be a 4 to equal 8.
If your first roll is a 5, your next roll would need to be a 3 to equal 8.
If your first roll is a 6, your next roll would need to be a 2 to equal 8, or a 6 to equal 12.


These are the only possibilities. There are 9 of them. Remember that number.


• 3. Now find the total amount of all possible combinations. If you learned the basics of combinations and permutations in school, which you should have at one point, then you should know that to find the total of all possible combinations from two unrelated events (without any repeating elements within an event) you simply multiply the total possibilities from each event together. There are six sides on each of the dice.


6 x 6 = 36

If you didn't know that and you've at least made it to Alg II, then may the math gods smite both you and your Alg teachers because you should have damn well covered it by now. Hell, I learned this basic crap in elementary school.

...

Sorry, thinking about the school system not doing it's job makes me kinda swear-y.

• 4. Your winning chance is the number of winning possibilities over the number of total possibilities. In other words, 9 over 36, or, to simplify, 1/4.

So there is a 1/4 chance you will win $8. Since there are not other possibilities except losing $3, then you have a 3/4 chance of losing $3.

1/4 + 3/4 = 1

25% + 75% = 100%

There's no way to make my logic any clearer than this.

(Oh, and make sure that all your fractions have the same denominator for the next step.)

• 5. So now you know the chances for each. Next, to find the expected win/loss per play, multiply the numerator of each fraction to the corresponding money value (using negative numbers to represent the loss), add them together, and divide by the denominator.

 ( 1 x $8 ) + ( 3 x -$3 ) = $8 + (-$9) = -$1
-$1 / 4 = -$0.25

In other words, if you play this game with your friend, over time you will average to losing 25 cents per round.

• Answer 1: Tell your friend to piss off because you're not an idiot. If he were really your friend he wouldn't try to take your money, or at least he'd check his damn math first so that the game would be fair for the two of you, because you are going to LOSE 25 cents per round!

Friends don't let friends do math incorrectly.

Speaking of which, I will be cutting off this lesson for now, but in part two I will be covering how to design a dice game so that you can make it as fair (or unfair) as you'd like.

• O • O • O • O • O •

Less than three,
The Artist